# Atomic sheaves on hyper-Kähler manifolds via Bridgeland moduli spaces

### v2

@@ 6.1.1. GM threefolds in a fixed GM fourfold
constructed in [49] and [15]. As explained in [36, Section 5.1] (see also [22]), we have a commutative diagram
@@
where the left column is the natural inclusion and the right column is the finite and unramified morphism in Lemma 6.2(2). Using this relation, the following result is obtained in [36, Proposition 5.2].
@@
Since $X$ is very general, by [25, Corollary B.9], there is a commutative diagram
@@
As $X$ and $Y$ are general, we know that $F(X)$ and $F(Y)$ are both smooth by [36, Theorem 3.2] and [49]. Moreover, from [17, Theorem B.2], we may assume that
@@
then $\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}$ and $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ are smooth by [19, Theorem 5.2]. Consequently, according to Proposition 6.4, $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is an immersed Lagrangian submanifold.
Recall that the moduli stacks of ordinary GM varieties of dimension $n=3$ or $4$ are smooth irreducible separated Deligne–Mumford stacks of finite type over $\mathbb{C}$ (cf. [21, Corollary 5.12] and [41, Proposition A.2]). Since the locus of Hodge-special GM fourfolds is the union of countably many divisors by [16, Corollary 4.6], using Lemma 6.9, we can construct a smooth family of ordinary GM fourfolds $\mathcal{X}\to C$ over a smooth connected curve $C$ and a smooth divisor $\mathcal{Y}\subset\mathcal{X}$ such that the fibers over a point $0\in C$ are $\mathcal{X}_{0}\cong X$ and $\mathcal{Y}_{0}\cong Y$. Moreover, there exists an uncountable set $F\subset C$ such that $\mathcal{X}_{c}$ is very general in the sense of Lemma 6.6 for each $c\in F$. Indeed, we start by considering a curve $C^{\prime}$ along with a family $\mathcal{X}^{\prime}\to C^{\prime}$ that satisfies the above assumptions. Next, we lift $C^{\prime}$ to a curve $C$ in the relative Hilbert scheme of hyperplane sections of $\mathcal{X}^{\prime}\to C^{\prime}$, obtaining the family $\mathcal{Y}\subset\mathcal{X}\to C$.
@@
Applying [21, Theorem 3.7, Proposition 4.1] to $\mathcal{X}\to C$ and $\mathcal{Y}\to C$, we obtain two families of Lagrangian data $(C,\mathscr{V}_{6},\mathscr{V}_{5},\mathscr{A}(\mathcal{X}))$ and $(C,\mathscr{V}_{6},\mathscr{V}_{5},\mathscr{A}(\mathcal{Y}))$ over $C$ [^7] See [21, Definition 3.15] for the precise definition., where $\mathscr{V}_{6}$, $\mathscr{V}_{5}$, $\mathscr{A}(\mathcal{X})$, and $\mathscr{A}(\mathcal{Y})$ are all locally free sheaves on $C$ such that
@@
for each $c\in C$. After replacing $C$ with an open dense neighborhood of $0\in C$, we can assume that these bundles are all trivial. Therefore, as in the proofs of [19, Theorem 5.2] and [18, Proposition 5.27], we get a family $p\colon\widetilde{\mathscr{Y}}_{\mathcal{X}}\to C$ of double dual EPW sextics and a family $q\colon\widetilde{\mathscr{Y}}^{\geq 2}_{\mathcal{Y}}\to C$ of double dual EPW surfaces such that $p^{-1}(c)\cong\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{c})^{\perp}}$ and $q^{-1}(c)\cong\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{c})^{\perp}}^{\geq 2}$ for each $c\in C$ by replacing $C$ with an open dense neighborhood of $0\in C$. Note that by our assumptions of families $\mathcal{X}$ and $\mathcal{Y}$, both $p$ and $q$ are smooth. We denote by $\mathscr{Y}_{\mathcal{X}}\to C$, $\mathscr{Y}_{\mathcal{Y}}^{\geq 2}\to C$, and $\mathscr{Y}^{\geq 2}_{\mathcal{X}}\to C$ the corresponding families of dual EPW sextics and dual EPW surfaces, respectively. Then $\mathscr{Y}_{\mathcal{X}},\mathscr{Y}^{\geq 2}_{\mathcal{X}}$ and $\mathscr{Y}^{\geq 2}_{\mathcal{Y}}$ are smooth. Note that they are all closed subschemes of $\mathbb{P}(\mathscr{V}^{\vee}_{6})\cong\mathbb{P}(V^{\vee}_{6})\times C$ and $\mathscr{Y}^{\geq 2}_{\mathcal{Y}_{c}}\subset\mathscr{Y}_{\mathcal{X}_{c}}$ for each $c\in C$, we see $\mathscr{Y}^{\geq 2}_{\mathcal{Y}}\subset\mathscr{Y}_{\mathcal{X}}$. Therefore, by passing to double coverings, taking simultaneous normalization (cf. [39, Theorem 12]), and using a relative version of Lemma 6.2, we have a commutative diagram
@@
Up to now, we have shrunk $C$ near $0\in C$ finitely many times. As $F$ is an infinite set, there is a point $t\in C\cap F$ such that Lemma 6.6 can be applied to $\mathcal{Y}_{t}\hookrightarrow\mathcal{X}_{t}$, i.e. $j_{t}\colon\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{t})^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{t})^{\perp}}$ is an immersed atomic Lagrangian submanifold. Therefore, using Proposition 4.4, we know that the Lagrangian submanifold $i_{c}\colon\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{c})^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{c})^{\perp}}$ is atomic for each $c\in C$, and in particular, $j_{0}=i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is atomic. Finally, the computation of the cohomology group is done in [20, Proposition 2.5] and [49, Proposition 0.5], and the orthogonal complement of the kernel of the restriction map is computed in Corollary 5.2 and Remark 3.11. ∎
@@
To this end, we first assume that $X$ is projective. Let $l\colon X\hookrightarrow\mathbb{P}^{n}$ be an embedding, then each hyperplane $H\subset\mathbb{P}^{n}$ intersects with each divisor in $\{D_{i}\}_{i\in\mathbb{N}}$. By [45, Theorem 3.3.1], for each $t\in\mathbb{N}$, there exists an open dense subset $U_{t}\subset|\operatorname{\mathcal{O}}_{\mathbb{P}^{n}}(1)|$ such that when $[H]\in U_{t}$, $H\cap X$ and $H\cap D_{i}$ are irreducible and $H$ intersects with $D_{i}$ and $D_{i}\cap D_{j}$ properly for each $i,j\leq t,i\neq j$. Thus, $\dim H\cap D_{i}\cap D_{j}\leq\dim X-3$ for $i,j\leq t,i\neq j$. We define $U:=\cap_{t\in\mathbb{N}}U_{t}$. Note that $U$ contains a closed point in $|\operatorname{\mathcal{O}}_{\mathbb{P}^{n}}(1)|$ as it is the complement of the union of countably many closed subvarieties and we work over $\mathbb{C}$. Then any $[H]\in U$ satisfies the required conditions. For the general case, let $\overline{X}$ be a projective variety containing $X$ as an open dense subset. Then $\overline{X}\setminus X$ has codimension at least one, and we can assume that there exist irreducible divisors $E_{j}\subset\overline{X}$ for $1\leq j\leq m$ such that $\overline{X}\setminus X\subset\cup_{1\leq j\leq m}E_{j}$. Thus, applying the above result to $\overline{X}$ and $\{E_{j}\}_{1\leq j\leq m}\cup\{D_{i}\}_{i\in\mathbb{N}}$, we get an irreducible curve $\overline{C}\subset\overline{X}$. Then $C:=\overline{C}\cap X$ satisfies all desired properties in the statement. ∎
@@ 6.1.2. GM fivefolds containing a fixed GM fourfold
According to [36, Section 5.2] and [22], there is a $\mathbb{P}^{1}$-fibration $\pi_{Y}\colon F_{qs}(Y)\twoheadrightarrow\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ and we have a commutative diagram
@@
where $F_{qs}(Y)$ is a connected component of the Hilbert scheme of quadric surfaces in $Y$ such that $S\cap X$ is a conic for each $[S]\in F_{qs}(Y)$ [^8] A quadric surface $S\subset Y$ is in $F_{qs}(Y)$ if and only if $S\subset\operatorname{Gr}(2,V_{4})$ for a $4$-dimensional subspace $V_{4}\subset V_{5}$. It is called a $\tau$-quadric in [16, Section 7.3].. Using this relation, the following result is obtained in [36, Proposition 5.6], which is an analog of Proposition 6.4.
@@
is of rank one. And from [19, Theorem 5.2(2)], we know that
@@
We can also demonstrate that $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is not injective as follows. If $i$ is an embedding, then by Theorem 6.7 and 6.13, $\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\subset\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is an atomic embedded Lagrangian. However, by [54, Equation (7.11)], we have
@@
which contradicts $\chi(\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2})=2\chi(\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2})=384$ by [24, Proposition 1.10].
@@ 6.2. Double EPW cubes
is constructed in [19, Theorem 5.7], such that $\widetilde{\mathsf{Z}}^{\geq 3}_{A(X)}$ is a smooth connected projective threefold when $X$ is general.
@@
Moreover, in [25, Section 8], the authors construct a rational map
@@
where $\pi_{Y}$ is induced by the projection functor, and $r$ is the morphism in Theorem 3.7. Moreover, $\pi_{Y}$ maps birationally onto a connected component of $M_{\sigma_{Y}}^{Y}(1,-1)$ by [25, Theorem C.13], and $r$ is an embedding when restricted to this component by [25, Lemma 8.20].
In particular, denoting by $L_{Y}$ the image of $r\circ\pi_{Y}$, it follows that $L_{X}$ is isomorphic to a connected component of $M_{\sigma_{Y}}^{Y}(1,-1)$ and $L_{Y}\subset M^{X}_{\sigma_{X}}(1,-1)\cong\widetilde{\mathsf{Z}}^{\geq 2}_{A(X)}$ is a Lagrangian subvariety. Moreover, $L_{Y}$ is smooth by [25, Theorem C.13, Corollary C.14] and $\mathrm{Hilb}^{3t+1}_{Y}\twoheadrightarrow L_{Y}$ is a divisorial contraction. Thus, Theorem 5.8 can be applied in this case, yielding:
@@ 7.1. The relation to the relative Jacobian fibration of GM fivefolds
First, we show that $s$ is injective, which is equivalent to showing that for any two different points $[Y],[Y^{\prime}]\in\mathcal{B}_{X}$, we have $\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}\neq\overline{\mathsf{Y}}^{\geq 2}_{A(Y^{\prime})^{\perp}}$. Indeed, if $\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}=\overline{\mathsf{Y}}^{\geq 2}_{A(Y^{\prime})^{\perp}}$, then we have $\mathsf{Y}^{\geq 2}_{A(Y)^{\perp}}=\mathsf{Y}^{\geq 2}_{A(Y^{\prime})^{\perp}}\subset\mathsf{Y}_{A(X)^{\perp}}$. However, by [20, Proposition 2.5], we have a canonical identification $A(Y)=A(Y^{\prime})$, which implies that $Y=Y^{\prime}$ by [17, Propoition 3.14].
@@
Note that $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are reflexive sheaves by [30, Corollary 1.2]. Let $U\subset\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ be the smooth locus and $V\subset\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ be the pullback of $U$. Hence, $V\cong U$ and $\mathcal{N}_{1}|_{V}\cong\mathcal{N}_{2}|_{U}$. From the fact that $\dim(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}\setminus V)=0$ by Lemma 6.2 and $\mathcal{N}_{2}$ is reflexive, we have
@@
Since $\mathcal{H}$ is smooth along $s(\mathcal{B}_{X})$, according to [23, Theorem 8.1], $\operatorname{Pic}^{0}(\widetilde{\mathscr{Y}}^{\geq 2}/\mathcal{B}_{X})$ is equipped with a holomorphic symplectic two-form and the relative Picard fibration
@@
is a Lagrangian fibration. In fact, while [23, Theorem 8.1] is stated only for embedded Lagrangians, it can also be applied in our case by Lemma 7.1. This is because the proof of [23, Theorem 8.1] involves a local study, as explained in [35, Section 3.1].
We denote by $\mathcal{M}$ the component of stable sheaves on $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ containing $i_{*}\mathcal{L}$, where $[Y]\in\mathcal{B}_{X}$ and $\mathcal{L}\in\operatorname{Pic}^{0}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})$. Note that $i_{*}\mathcal{L}$ is stable because it is a rank one torsion-free sheaf supported on the image of the normalization map $i$. Then $\mathcal{M}$ has an open dense subscheme $\mathcal{M}^{0}$ isomorphic to $\operatorname{Pic}^{0}(\widetilde{\mathscr{Y}}^{\geq 2}/\mathcal{B}_{X})$ such that the support morphism $\mathcal{M}^{0}\to\mathcal{B}_{X}$ is isomorphic to the relative Picard fibration via the identification of $\mathcal{B}_{X}$ as an open subscheme of $\mathcal{H}$ by Lemma 7.1. Thus, $\mathcal{M}^{0}$ also carries a holomorphic symplectic structure. This strategy is explained in [54, Section 7.4].
On the other hand, according to [20, Theorem 1.1], there is a canonical isomorphism of principally polarized abelian varieties
@@
for any $[Y]\in\mathcal{B}_{X}$. Thus, $\mathrm{J}(Y)\cong\mathrm{Alb}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})\cong\operatorname{Pic}^{0}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})$ and the fiber of $\mathcal{M}^{0}\to\mathcal{B}_{X}$ over $[Y]\in\mathcal{B}_{X}$ is canonically isomorphic to $\mathrm{J}(Y)$. This situation is very similar to the case of cubic fourfolds (cf. [23, Example 8.22] and [51, 35]), and we hope that the techniques in [51] can be applied to the Gushel–Mukai case to construct a suitable compactification $\overline{\mathcal{M}}$ of $\mathcal{M}^{0}\to\mathcal{B}_{X}$.

### v3

@@ 6.1.1. GM threefolds in a fixed GM fourfold
constructed in [54] and [16]. As explained in [40, Section 5.1] (see also [23]), we have a commutative diagram
@@
where the left column is the natural inclusion and the right column is the finite and unramified morphism in Lemma 6.2(2). Using this relation, the following result is obtained in [40, Proposition 5.2].
@@
Since $X$ is very general, by [27, Corollary B.9], there is a commutative diagram
@@
As $X$ and $Y$ are general, we know that $F(X)$ and $F(Y)$ are both smooth by [40, Theorem 3.2] and [54]. Moreover, from [18, Theorem B.2], we may assume that
@@
then $\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}$ and $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ are smooth by [20, Theorem 5.2]. Consequently, according to Proposition 6.4, $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is an immersed Lagrangian submanifold.
Recall that the moduli stacks of ordinary GM varieties of dimension $n=3$ or $4$ are smooth irreducible separated Deligne–Mumford stacks of finite type over $\mathbb{C}$ (cf. [22, Corollary 5.12] and [45, Proposition A.2]). Since the locus of Hodge-special GM fourfolds is the union of countably many divisors by [17, Corollary 4.6], using Lemma 6.9, we can construct a smooth family of ordinary GM fourfolds $\mathcal{X}\to C$ over a smooth connected curve $C$ and a smooth divisor $\mathcal{Y}\subset\mathcal{X}$ such that the fibers over a point $0\in C$ are $\mathcal{X}_{0}\cong X$ and $\mathcal{Y}_{0}\cong Y$. Moreover, there exists an uncountable set $F\subset C$ such that $\mathcal{X}_{c}$ is very general in the sense of Lemma 6.6 for each $c\in F$. Indeed, we start by considering a curve $C^{\prime}$ along with a family $\mathcal{X}^{\prime}\to C^{\prime}$ that satisfies the above assumptions. Next, we lift $C^{\prime}$ to a curve $C$ in the relative Hilbert scheme of hyperplane sections of $\mathcal{X}^{\prime}\to C^{\prime}$, obtaining the family $\mathcal{Y}\subset\mathcal{X}\to C$.
@@
Applying [22, Theorem 3.7, Proposition 4.1] to $\mathcal{X}\to C$ and $\mathcal{Y}\to C$, we obtain two families of Lagrangian data $(C,\mathscr{V}_{6},\mathscr{V}_{5},\mathscr{A}(\mathcal{X}))$ and $(C,\mathscr{V}_{6},\mathscr{V}_{5},\mathscr{A}(\mathcal{Y}))$ over $C$ [^9] See [22, Definition 3.15] for the precise definition., where $\mathscr{V}_{6}$, $\mathscr{V}_{5}$, $\mathscr{A}(\mathcal{X})$, and $\mathscr{A}(\mathcal{Y})$ are all locally free sheaves on $C$ such that
@@
for each $c\in C$. After replacing $C$ with an open dense neighborhood of $0\in C$, we can assume that these bundles are all trivial. Therefore, as in the proofs of [20, Theorem 5.2] and [19, Proposition 5.27], we get a family $p\colon\widetilde{\mathscr{Y}}_{\mathcal{X}}\to C$ of double dual EPW sextics and a family $q\colon\widetilde{\mathscr{Y}}^{\geq 2}_{\mathcal{Y}}\to C$ of double dual EPW surfaces such that $p^{-1}(c)\cong\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{c})^{\perp}}$ and $q^{-1}(c)\cong\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{c})^{\perp}}^{\geq 2}$ for each $c\in C$ by replacing $C$ with an open dense neighborhood of $0\in C$. Note that by our assumptions of families $\mathcal{X}$ and $\mathcal{Y}$, both $p$ and $q$ are smooth. We denote by $\mathscr{Y}_{\mathcal{X}}\to C$, $\mathscr{Y}_{\mathcal{Y}}^{\geq 2}\to C$, and $\mathscr{Y}^{\geq 2}_{\mathcal{X}}\to C$ the corresponding families of dual EPW sextics and dual EPW surfaces, respectively. Then $\mathscr{Y}_{\mathcal{X}},\mathscr{Y}^{\geq 2}_{\mathcal{X}}$ and $\mathscr{Y}^{\geq 2}_{\mathcal{Y}}$ are smooth. Note that they are all closed subschemes of $\mathbb{P}(\mathscr{V}^{\vee}_{6})\cong\mathbb{P}(V^{\vee}_{6})\times C$ and $\mathscr{Y}^{\geq 2}_{\mathcal{Y}_{c}}\subset\mathscr{Y}_{\mathcal{X}_{c}}$ for each $c\in C$, we see $\mathscr{Y}^{\geq 2}_{\mathcal{Y}}\subset\mathscr{Y}_{\mathcal{X}}$. Therefore, by passing to double coverings, taking simultaneous normalization (cf. [43, Theorem 12]), and using a relative version of Lemma 6.2, we have a commutative diagram
@@
Up to now, we have shrunk $C$ near $0\in C$ finitely many times. As $F$ is an infinite set, there is a point $t\in C\cap F$ such that Lemma 6.6 can be applied to $\mathcal{Y}_{t}\hookrightarrow\mathcal{X}_{t}$, i.e. $j_{t}\colon\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{t})^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{t})^{\perp}}$ is an immersed atomic Lagrangian submanifold. Therefore, using Proposition 4.4, we know that the Lagrangian submanifold $i_{c}\colon\widetilde{\mathsf{Y}}_{A(\mathcal{Y}_{c})^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(\mathcal{X}_{c})^{\perp}}$ is atomic for each $c\in C$, and in particular, $j_{0}=i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is atomic. Finally, the computation of the cohomology group is done in [21, Proposition 2.5] and [54, Proposition 0.5], and the orthogonal complement of the kernel of the restriction map is computed in Corollary 5.2 and Remark 3.11. ∎
@@
To this end, we first assume that $X$ is projective. Let $l\colon X\hookrightarrow\mathbb{P}^{n}$ be an embedding, then each hyperplane $H\subset\mathbb{P}^{n}$ intersects with each divisor in $\{D_{i}\}_{i\in\mathbb{N}}$. By [49, Theorem 3.3.1], for each $t\in\mathbb{N}$, there exists an open dense subset $U_{t}\subset|\operatorname{\mathcal{O}}_{\mathbb{P}^{n}}(1)|$ such that when $[H]\in U_{t}$, $H\cap X$ and $H\cap D_{i}$ are irreducible and $H$ intersects with $D_{i}$ and $D_{i}\cap D_{j}$ properly for each $i,j\leq t,i\neq j$. Thus, $\dim H\cap D_{i}\cap D_{j}\leq\dim X-3$ for $i,j\leq t,i\neq j$. We define $U:=\cap_{t\in\mathbb{N}}U_{t}$. Note that $U$ contains a closed point in $|\operatorname{\mathcal{O}}_{\mathbb{P}^{n}}(1)|$ as it is the complement of the union of countably many closed subvarieties and we work over $\mathbb{C}$. Then any $[H]\in U$ satisfies the required conditions. For the general case, let $\overline{X}$ be a projective variety containing $X$ as an open dense subset. Then $\overline{X}\setminus X$ has codimension at least one, and we can assume that there exist irreducible divisors $E_{j}\subset\overline{X}$ for $1\leq j\leq m$ such that $\overline{X}\setminus X\subset\cup_{1\leq j\leq m}E_{j}$. Thus, applying the above result to $\overline{X}$ and $\{E_{j}\}_{1\leq j\leq m}\cup\{D_{i}\}_{i\in\mathbb{N}}$, we get an irreducible curve $\overline{C}\subset\overline{X}$. Then $C:=\overline{C}\cap X$ satisfies all desired properties in the statement. ∎
@@ 6.1.2. GM fivefolds containing a fixed GM fourfold
According to [40, Section 5.2] and [23], there is a $\mathbb{P}^{1}$-fibration $\pi_{Y}\colon F_{qs}(Y)\twoheadrightarrow\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ and we have a commutative diagram
@@
where $F_{qs}(Y)$ is a connected component of the Hilbert scheme of quadric surfaces in $Y$ such that $S\cap X$ is a conic for each $[S]\in F_{qs}(Y)$ [^10] A quadric surface $S\subset Y$ is in $F_{qs}(Y)$ if and only if $S\subset\operatorname{Gr}(2,V_{4})$ for a $4$-dimensional subspace $V_{4}\subset V_{5}$. It is called a $\tau$-quadric in [17, Section 7.3].. Using this relation, the following result is obtained in [40, Proposition 5.6], which is an analog of Proposition 6.4.
@@
is of rank one. And from [20, Theorem 5.2(2)], we know that
@@
We can also demonstrate that $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is not injective as follows. If $i$ is an embedding, then by Theorem 6.7 and 6.13, $\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\subset\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is an atomic embedded Lagrangian. However, by [59, Equation (7.11)], we have
@@
which contradicts $\chi(\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2})=2\chi(\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2})=384$ by [25, Proposition 1.10].
@@ 6.2. Double EPW cubes
is constructed in [20, Theorem 5.7], such that $\widetilde{\mathsf{Z}}^{\geq 3}_{A(X)}$ is a smooth connected projective threefold when $X$ is general.
@@
Moreover, in [27, Section 8], the authors construct a rational map
@@
where $\pi_{Y}$ is induced by the projection functor, and $r$ is the morphism in Theorem 3.7. Moreover, $\pi_{Y}$ maps birationally onto a connected component of $M_{\sigma_{Y}}^{Y}(1,-1)$ by [27, Theorem C.13], and $r$ is an embedding when restricted to this component by [27, Lemma 8.20].
In particular, denoting by $L_{Y}$ the image of $r\circ\pi_{Y}$, it follows that $L_{X}$ is isomorphic to a connected component of $M_{\sigma_{Y}}^{Y}(1,-1)$ and $L_{Y}\subset M^{X}_{\sigma_{X}}(1,-1)\cong\widetilde{\mathsf{Z}}^{\geq 2}_{A(X)}$ is a Lagrangian subvariety. Moreover, $L_{Y}$ is smooth by [27, Theorem C.13, Corollary C.14] and $\mathrm{Hilb}^{3t+1}_{Y}\twoheadrightarrow L_{Y}$ is a divisorial contraction. Thus, Theorem 5.8 can be applied in this case, yielding:
@@ 7.1. The relation to the relative Jacobian fibration of GM fivefolds
First, we show that $s$ is injective, which is equivalent to showing that for any two different points $[Y],[Y^{\prime}]\in\mathcal{B}_{X}$, we have $\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}\neq\overline{\mathsf{Y}}^{\geq 2}_{A(Y^{\prime})^{\perp}}$. Indeed, if $\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}=\overline{\mathsf{Y}}^{\geq 2}_{A(Y^{\prime})^{\perp}}$, then we have $\mathsf{Y}^{\geq 2}_{A(Y)^{\perp}}=\mathsf{Y}^{\geq 2}_{A(Y^{\prime})^{\perp}}\subset\mathsf{Y}_{A(X)^{\perp}}$. However, by [21, Proposition 2.5], we have a canonical identification $A(Y)=A(Y^{\prime})$, which implies that $Y=Y^{\prime}$ by [18, Propoition 3.14].
@@
Note that $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are reflexive sheaves by [33, Corollary 1.2]. Let $U\subset\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ be the smooth locus and $V\subset\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ be the pullback of $U$. Hence, $V\cong U$ and $\mathcal{N}_{1}|_{V}\cong\mathcal{N}_{2}|_{U}$. From the fact that $\dim(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}\setminus V)=0$ by Lemma 6.2 and $\mathcal{N}_{2}$ is reflexive, we have
@@
Proof of Corollary 1.4.
This follows from Theorem 6.13 and the same argument as in Theorem 1.2. Note that the dimension of the deformation space is $20$ since the morphism $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is a closed embedding away from finitely many points by Lemma 6.2(2) and there exists a nice family of immersed Lagrangians of $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ as required in Lemma 5.12. Indeed, the family $L_{V}\to V$ in Lemma 5.12 is constructed by a relative version of Theorem 6.13 as in the proof of Theorem 6.7, which satisfies (1) and (2) in Lemma 5.12 by Lemma 6.2. The assumption (3) in Lemma 5.12 follows from Lemma 7.1. Hence, we can apply Lemma 5.12 in this case to determine the dimension of the deformation space. ∎
Since $\mathcal{H}$ is smooth along $s(\mathcal{B}_{X})$, according to [24, Theorem 8.1], $\operatorname{Pic}^{0}(\widetilde{\mathscr{Y}}^{\geq 2}/\mathcal{B}_{X})$ is equipped with a holomorphic symplectic two-form and the relative Picard fibration
@@
is a Lagrangian fibration. In fact, while [24, Theorem 8.1] is stated only for embedded Lagrangians, it can also be applied in our case by Lemma 7.1. This is because the proof of [24, Theorem 8.1] involves a local study, as explained in [39, Section 3.1].
We denote by $\mathcal{M}$ the component of stable sheaves on $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ containing $i_{*}\mathcal{L}$, where $[Y]\in\mathcal{B}_{X}$ and $\mathcal{L}\in\operatorname{Pic}^{0}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})$. Note that $i_{*}\mathcal{L}$ is stable because it is a rank one torsion-free sheaf supported on the image of the normalization map $i$. Then $\mathcal{M}$ has an open dense subscheme $\mathcal{M}^{0}$ isomorphic to $\operatorname{Pic}^{0}(\widetilde{\mathscr{Y}}^{\geq 2}/\mathcal{B}_{X})$ such that the support morphism $\mathcal{M}^{0}\to\mathcal{B}_{X}$ is isomorphic to the relative Picard fibration via the identification of $\mathcal{B}_{X}$ as an open subscheme of $\mathcal{H}$ by Lemma 7.1. Thus, $\mathcal{M}^{0}$ also carries a holomorphic symplectic structure. This strategy is explained in [59, Section 7.4].
On the other hand, according to [21, Theorem 1.1], there is a canonical isomorphism of principally polarized abelian varieties
@@
for any $[Y]\in\mathcal{B}_{X}$. Thus, $\mathrm{J}(Y)\cong\mathrm{Alb}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})\cong\operatorname{Pic}^{0}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})$ and the fiber of $\mathcal{M}^{0}\to\mathcal{B}_{X}$ over $[Y]\in\mathcal{B}_{X}$ is canonically isomorphic to $\mathrm{J}(Y)$. This situation is very similar to the case of cubic fourfolds (cf. [24, Example 8.22] and [56, 39]), and we hope that the techniques in [56] can be applied to the Gushel–Mukai case to construct a suitable compactification $\overline{\mathcal{M}}$ of $\mathcal{M}^{0}\to\mathcal{B}_{X}$.